OpenStudy (anonymous):

how do i simplify (8a)^2

OpenStudy (anonymous):

(8a)^2 = (8^2)(a^2)

OpenStudy (australopithecus):

(8a)^(2) = (8a)(8a)

OpenStudy (anonymous):

Please explain

OpenStudy (australopithecus):

\[x^2 = (x)(x)\] \[x^3 = (x)(x)(x)\] etc thus \[(xa)^2 = (xa)(xa)\]

OpenStudy (anonymous):

Ok Thanks

OpenStudy (australopithecus):

do you follow now?

OpenStudy (anonymous):

(8a)^-2

OpenStudy (australopithecus):

how did you get that

OpenStudy (anonymous):

Thats the actually question simplify (8a)^-2

OpenStudy (australopithecus):

with exponents you combine like terms \[(2x)^3 = (2)(2)(2)(x)(x)(x) = 4(2)(x)(x)(x) = 8(x)(x)(x) = 8x^3\]

OpenStudy (anonymous):

Ok I got it

OpenStudy (australopithecus):

\[(8a)^{2} \neq (8a)^{-2}\] \[(8a)^{-2} = \frac{1}{(8a)^{2}}\]

OpenStudy (australopithecus):

just a note so you understand negative exponents

OpenStudy (australopithecus):

\[x^{-1} = \frac{1}{x^{1}} = \frac{1}{x}\]